MENTORING PAIRINGS

APPRENTICE and SIT-INS Graduate students and e-mails

Name (weeks)  @math.uchicago.edu
Alan Anders  aanders
Khalid Bou-Rabee  khalid
Michael Geline (1-4)  geline
Bert Guillou (1-4)  bertg
Gabriel Kerr (1-6)  gdkerr
Aaron Marcus (one mentoree)  amarcus
Courtney Morris (1-4 or 6)  clmorris
Andrew Putman (1-4)  andyp
Ian Shipman (one mentoree)  ics
Saravanan Thiyagarajan  avan

Coordinator: Michael Geline Undergraduate students, e-mails, and mentors.

Name (32)  @uchicago.edu  mentor(s)
William Cole Abram  abramwc  Guillou, Morris
Dylan Allegretti  allegretti  Anders, Geline
Adam Anderson  adamanderson Kerr, Thiyagarajan
Elan Bechor  ebechor  Kerr, Thiyagarajan
Michael Calderbank  Mcalderbank  Guillou, Morris
Daniel Cervone  dcervone  Kerr, Thiyagarajan
Beau Dabbs  bdabbs13  Anders, Geline
Catherine Easton  catherine  Guillou, Morris
Mufti Nomaan Farooqui  mnomy88  Bou-Rabee, Putman
Hannah Gordon  hannahg  Bou-Rabee, Putman
Austin Gross  Agross  Bou-Rabee, Putman
Eric Daniel Jaffe  ericjaffe  Anders, Geline
Ethan Jerzak  jerzak  Guillou, Morris
Michael (Ryan) Julian  mrjulian  Kerr, Thiyagarajan
Katayun J. Kamdin  kkamdin  Kerr, Thiyagarajan
Suyeon Khim  skhim  Marcus, Shipman
Satory Kudose  skudose  Anders, Geline
William Leeb  weleeb  Kerr, Thiyagarajan
Chenchuan Li  chenchuan Bou-Rabee, Putman
Cong Han Lim  conghan  Kerr, Thiyagarajan
Ellen McCullagh  ellenm  Bou-Rabee, Putman
Scott Messick  smessick  Anders, Geline
Jamie Morgenstern  jamiemmt  Guillou, Morris
Matt Olson  molson2  Anders, Geline
Bo Peng  bopeng  Guillou, Morris
Andrew Poppick  apoppick  Anders, Geline
James Salvatore  jsal13  Guillou, Morris
Honi Sanders  honiraphael  Anders, Geline
Brad Sugarman  bvs  Bou-Rabee, Putman
Michael Travis  mtravis  Bou-Rabee, Putman
Brooke Ullery  bullery  Guillou, Morris
Gideon Tan Guang Yuan  serapuff  Kerr, Thiyagarajan
Tengren Zhang  tengren85  Bou-Rabee, Putman

SESAME Graduate students and e-mails

Name (weeks)  @math.uchicago.edu
Marius Beceanu (2-5)  mbeceanu
Anna-Marie Bohmann  bohmann
Tom Church  tchurch
Matthew Day  mattday
Shawn Drenning (1-7)  shawn
Damir Dzhafarov (2-8)  damir
Shawn Henry  sjhenry@umich.edu
Niles Johnson  niles
Irene Peng  irine
Travis Schedler  trasched

(Additional mentor(s) starting fifth week) Coordinator: Niles Johnson Undergraduate students, e-mails, and mentors.

Name (22)  @uchicago.edu  mentor(s)
Stephan Balady  sbalady  Beceanu, Peng
Sarah Bennett  sbennett  Bohmann, Johnson
David Berger  iloveyou  Beceanu, Peng
Chad Casarotto  chadmbol  Drenning, Schedler
Christopher Logan Chariker  chariker  Boyarchenko, Schedler
Le Chen  lelechen  Beceanu, Peng
David Coley  dcoley  Boyarchenko
Andrew Eisenberg  aeisen  Church, Day
Arpit Gupta  arpit  Drenning, Schedler
Adam Kaye  adamkaye  Church, Day
R. Andrew Kreek  rakreek  Church, Day
Joshua Blaise Leners  leners  Dzhafarov, Henry
Carmel Levy  carmel1  Dzhafarov, Henry
Yana Peysakhovich  yana  Drenning, Schedler
Robert Prag  rprag  Bohmann, Johnson
Edward Ruggeri  ruggeri  Bohmann, Johnson
Maxwell Shron  mshron  Church, Day
Andries Smith  andries  Bohmann, Johnson
Michael Thvedt  cynicalk  Bohmann, Johnson
Alexander Volfovsky  alexvolf  Beceanu, Peng
Matthew Woolf  mwoolf  Church, Day
Qian Sophia Zhang  qian  Beceanu, Peng

YSP and COLLEGIATE SCHOLARS Graduate students and e-mails

Name  @math.uchicago.edu
Sundeep Balaji  sundeepb
Ian Biringer (1-2)  biringer
Mitya Boyarchenko  mitya
Michael Broshi (1-6, 8)  mbroshi
Spencer Dowdall (3-8)  sdowdall
Jim Fowler  fowler
Shawn Henry  sjhenry@umich.edu
John Lind  lind
Justin Malestein  justinm
Aaron Marcus  amarcus
Ian Shipman  ics

Coordinator: Jim Fowler

Undergraduate students, e-mails, and mentors.

YSP: 7–8

Name (8)  @uchicago.edu  mentor(s)
Emily Bargar  ebargar  Balaji, Broshi
Padraic James Bartlett  padraic  Lind, Malestein
Sarah Constantin  sconsta1  Balaji, Broshi
John Dyer  dyer  Balaji, Broshi
Gabriel Gaster  eulerian  Balaji, Broshi
Alice Mark  amark  Marcus, Shipman
James Marrone  jmar  Biringer-Dowdall, Fowler
Brian Taylor  bgtaylor  Balaji, Broshi

YSP: 9–10

Name (8)  @uchicago.edu  mentor(s)
Nicholas Ford  nford  Henry, Dzhafarov
Daniel Gardner  dbgardner  Boyarchenko
Drew A. Hudec  dhudec  Marcus, Shipman
Grant Larsen  gel  Marcus, Shipman
Jean Morrison  jeanm  Lind, Malestein
Alexander J. Munk  amunk  Lind, Malestein
Prerna Nadathur  pnadathur Biringer-Dowdall, Fowler
Caitlin Waddle kwaddle Lind, Malestein

YSP: 11–12

Name (6)  @uchicago.edu  mentor(s)
Harry Altman  haltman  Biringer-Dowdall, Fowler
Corrin Clarkson  clarkson  Biringer-Dowdall, Fowler
David Cohen  dcohen1  Marcus, Shipman
Lucas Culler  luccul  Boyarchenko
Jonathan Nieder  jrnieder  Boyarchnko
Sam Raskin  sraskin  Boyarchenko

COLLEGIATE SCHOLARS

Name (2)  @uchicago.edu  mentor(s)
Fan Fei Chong  ffclark  Lind, Malestein
Calvin Lin Zhiwei  linzw  Biringer-Dowdall, Fowler